# Token Embeddings Violate the Manifold Hypothesis

### v2

@@ Appendix A Supplementary Material
A.1. Background
@@
Figure 4. Data flow in a typical LLM. A sequence of tokens forming the query is converted via the token input embedding $e^{w}$ into the initial context window, as a point in the latent space $X^{w}$. Each of these windows in the latent space are converted, token-by-token, into probability distributions via $f$ into the single token latent space $X$. From these, each token presented in the output (in the set $Y$) is obtained via a random draw. These output tokens are then used for subsequent windows.
@@
Next window prediction:
@@
A.2. Computational resource requirements
Applying our entire method (starting with the token embedding matrix and ending with the $p$-values for each of the three tests at every token) to each model took approximately 12 hours of wall clock time (per model) on an Intel Core i7-3820 with 32 GB of CPU RAM and no GPU running at 3.60GHz. Almost all of the runtime was spent computing the token distance matrix. The subsequent stages, including the three hypothesis tests, shown in Figure 2 run in only a few minutes on the same machine. The distance matrix calculation was performed using scipy.spatial.distance_matrix(), which produces a dense matrix, by computing the distance between all pairs of tokens in the latent space. Because of the relatively low runtime requirements, we did not feel the need to use sparsity-enhancing techniques to accelerate the distance matrix calculation.
A.3. Interpretation of fiber bundles in terms of spatial variability
@@
Figure 5. Our method applied to a fiber bundle in $\mathbb{R}^{2}$. The vertical direction is the base space (long spatial scale), while the horizontal direction represents the fibers space (short spatial scale). Gray points on the right frame show estimates from a random sampling of points in the strip; the solid line shows the theoretical area versus radius curve.
@@
Figure 5 shows an example of this situation. It consists of a $1$-dimensional base space (the long spatial scale) and $1$-dimensional fibers (the short spatial scale), which in this case forms a narrow strip in the plane. Volumes (areas, in this case) of balls of small radius scale quadratically (slope $2$ in a log-log plot), but scale asymptotically linearly (slope $1$ in a log-log plot) for large radii. The transition between these two behaviors is detectable by way of a corner in the plot. This situation is easily and robustly estimated from the data; the gray points in Figure 5 (right) are derived from a random sampling of points drawn from the strip.
@@
To test for multiple spatial scales, we propose a signal model that is mathematically represented by a fiber bundle. In a fiber bundle, the long spatial scales are modeled by a space $S$, but all possible measurements are modeled by a function $p:E\to S$. The idea is that fibers $p^{-1}(b)$ are still cartesian products: pairs of both spatial scales, and these are all identical up to diffeomorphism. Our method relies upon a particular geometric property of fiber bundles: we can identify if the fibers are not all identical according to when the conclusion of Theorem 1 is violated.
@@
One possible way that a fiber bundle might arise is if the underlying noise model is homooscedastic, which is to say that the dimensionality of the noise near a token does not depend on that token. In this case, the fibers (short spatial scales) mostly correspond to noise, and the base (long spatial scales) correspond mostly to signal. Nevertheless, a fiber bundle may arise with noiseless data as well; the presence of multiple spatial scales is what matters.
@@
On the other hand, if the underlying noise model is heteroscedastic—it depends on the token in question—then the fiber bundle model should be rejected. Figure 3(a) shows a situation that is not a fiber bundle, since there is a change in the dimension of the fiber. In the upper portion of the figure, the fiber dimension is 00 while in the lower portion the fiber dimension is $1$. This is detectable by looking at the volume versus radius plots for two samples. While both samples show corners in their volume versus radius plots, Theorem 1 establishes that the slopes always decrease with increasing radius for a fiber bundle. This is violated for the sample marked $\psi_{1}$, so we conclude that the space is not a fiber bundle. On the other hand, because the sample marked $\psi_{2}$ does not exhibit this violation, it is important to note that if a sample yields data consistent with Theorem 1, we cannot conclude that the space is a fiber bundle.
A.4. Mathematical proof of Theorems 1 and 2
@@
By the submersion theorem ([48]), if the Jacobian matrix of a fibered manifold $p:T\to S$ at every point has rank equal to the dimension of $S$, then the preimages $p^{-1}(\psi)\subseteq T$ of each point $\psi\in T$ are all diffeomorphic to each other. These preimages form the fibers discussed in the earlier sections of the paper.
@@
(of Theorem 1) If $e$ is assumed to be a smooth embedding [^4] Smoothness is not a strong constraint here. Any continuous function (such as ReLU) can be approximated well enough by a smooth map., the image of $e$ is a manifold of dimension $\dim T$. The pushforward of a volume form is a contravariant functor, so this means that $e_{*}v$ is the volume form for a Riemannian metric on $e(T)$. Using this Riemannian metric on $e(T)$, then [49, Thm 3.1] implies that for every $\psi\in e(T)$, if $r\ll\tau$, then
(2) $(e_{*}v)\left(B_{r}(\psi)\right)=O\left(r^{\dim T}\right).$
@@
Since $e$ is assumed to be an embedding, by the tubular neighborhood theorem ([48]), it must be that $r_{2}<\tau$. Define
@@
We now address Theorem 2.
@@
(of Theorem 2) We need to manipulate the inequalities in the hypothesis to show that they satisfy [50, Thm. 1]. Once the hypothesis is satisfied, [50, Thm. 1] establishes that a generic set of transformers will embed $T^{w}$ into the output of the LLM. Notice that [50, Thm. 1] considers the response of the LLM to a single token, though this can be easily extended by enlarging the bounding manifold dimension. With this in mind, the hypothesis to be satisfied is that
@@
A.5. Additional views for Figure 1
@@
Figure 6. Additional rotational views of the PCA of the neighborhood of ember in Mistral7B shown in Figure 1. Each point shown is a token. Colors are $p$-values for the manifold hypothesis, with darker colors corresponding to lower likelihood that that point’s neighborhood is a manifold with low curvature.
See Fig. 6.
A.6. Table of fiber bundle rejections in all four LLMs
Table 3. Violations of the fiber bundle hypothesis
Model Token Slope change $p$-value Comment
GPT2 laughable Smaller radius $9\times 10^{-6}$ Must start a word
GPT2 nuance Smaller radius $2\times 10^{-4}$ Must start a word
GPT2 dt Smaller radius $2\times 10^{-4}$
GPT2 Mesh Smaller radius $2\times 10^{-4}$
GPT2 affect Smaller radius $3\times 10^{-4}$ Must start a word
GPT2 Thankfully Smaller radius $3\times 10^{-4}$
GPT2 swat Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 Malaysian Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 Palestinian Smaller radius $7\times 10^{-4}$ Must start a word
GPT2 wins Smaller radius $8\times 10^{-4}$ Must start a word
GPT2 hedon Smaller radius $9\times 10^{-4}$
GPT2 donor Smaller radius $9\times 10^{-4}$ Must start a word
GPT2 Xan Larger radius $3\times 10^{-8}$ Must start a word
GPT2 aunder Larger radius $2\times 10^{-4}$
GPT2 Dri Larger radius $2\times 10^{-4}$
GPT2 ney Larger radius $3\times 10^{-4}$
GPT2 rodu Larger radius $3\times 10^{-4}$
GPT2 Insert Larger radius $4\times 10^{-4}$
GPT2 Ying Larger radius $4\times 10^{-4}$ Must start a word
Llemma7B pax Larger radius $3\times 10^{-4}$
Mistral7B änge Smaller radius $8\times 10^{-4}$
Mistral7B H0 Larger radius $5\times 10^{-4}$
Mistral7B monitor Larger radius $8\times 10^{-5}$ Must start a word
@@ A.6. Table of fiber bundle rejections in all four LLMs
A.7. Tables of manifold rejections in all four LLMs
Table 4. Rejections of the manifold hypothesis for GPT2 (1 of 2)
Number Token $p$-value Comment
1 Xan 0.0000000280 Must start a word
2 chemist 0.0000000602
3 fry 0.00000564 Must start a word
4 laughable 0.00000807 Must start a word
5 Retro 0.0000108 Must start a word
6 laughable 0.0000156 Must start a word
7 Thankfully 0.0000339
8 intoler 0.0000514 Must start a word
9 assumes 0.0000533 Must start a word
10 handwritten 0.0000783 Must start a word
11 MY 0.0000811 Must start a word
12 nuance 0.000105 Must start a word
13 playground 0.000110 Must start a word
14 dt 0.000124
15 aunder 0.000132
16 Mesh 0.000135
17 Disney 0.000139
18 McCl 0.000155 Must start a word
19 Dri 0.000196
20 Destiny 0.000206 Must start a word
21 ney 0.000206
22 juvenile 0.000209 Must start a word
23 advent 0.000211 Must start a word
24 perpet 0.000213 Must start a word
25 otally 0.000223
26 affect 0.000226 Must start a word
27 efficient 0.000238
28 rodu 0.000242
29 Kl 0.000260 Must start a word
30 Thankfully 0.000278
31 duct 0.000284
32 frail 0.000292 Must start a word
33 alle 0.000304 Must start a word
34 Cha 0.000322 Must start a word
35 OO 0.000333
Table 5. Rejections of the manifold hypothesis for GPT2 (2 of 2)
Number Token $p$-value Comment
36 Insert 0.000336
37 Ying 0.000357 Must start a word
38 Georgia 0.000373
39 flo 0.000384
40 uddled 0.000413
41 subsequ 0.000421 Must start a word
42 lig 0.000439 Must start a word
43 Environmental 0.000456 Must start a word
44 Vladimir 0.000466 Must start a word
45 Defence 0.000510 Must start a word
46 swat 0.000558 Must start a word
47 registering 0.000564 Must start a word
48 Malaysian 0.000569 Must start a word
49 trough 0.000575 Must start a word
50 ears 0.000615 Must start a word
51 Meh 0.000637 Must start a word
52 Palestinian 0.000648 Must start a word
53 reckless 0.000654 Must start a word
54 Wax 0.000658 Must start a word
55 Dist 0.000673 Must start a word
56 Quad 0.000725 Must start a word
57 relations 0.000749
58 wins 0.000781 Must start a word
59 KER 0.000785
60 gist 0.000786 Must start a word
61 hedon 0.000834
62 MMA 0.000839 Must start a word
63 donor 0.000865 Must start a word
64 aleb 0.000873
65 Asked 0.000873
66 Isaiah 0.000940 Must start a word
67 Survey 0.000947 Must start a word
68 youngster 0.000987 Must start a word
Table 6. Rejections of the manifold hypothesis for Llemma7B
Number Token $p$-value Comment
1 clojure 0.00000000434
2 agu 0.000000846
3 ¡Cyrillic¿ 0.000000938 Must start a word
4 custom 0.00000163
5 kW 0.00000496 Must start a word

### v3

@@ Appendix A Supplementary Material
A.1 Additional discussion of the literature
Although there a number of mathematically distinct definitions of dimension, they coincide for manifolds. Intrinsic dimension of a point cloud has been widely studied in the literature Schweinhart [2020]. All of the papers that address practical estimation of intrinsic dimension from data do so under the assumption that the data lie on a manifold. For instance, Levina and Bickel [2004] provides maximum likelihood estimates, operating under the assumption that data “can be efficiently summarized in a space of a much lower dimension, such as a nonlinear manifold.” Moreover, Facco et al. [2017] provides a surprisingly effective approach using extremely small neighborhoods, a fact which underlies our use of our test on a “small radius neighborhood” in Section 5.2.
Even though we ultimately reject the manifold hypothesis, the token subspace has been studied under the assumption that it is a manifold. For instance Lee et al. [2025] observe that,“Language models often construct similar local representations as constructed from local linear embeddings.” We note that most of the tokens in a typical vocabulary fail to reject the manifold hypothesis, which means that without the deeper analysis we pursue, it is easy to overlook important differences. Furthermore, our Table 2 shows significant topological differences in the local representations between models.
As introduced in Section 1.2, we mentioned that Lee et al. [2025] show a few things that are consistent with our findings. They note that “Token embeddings exhibit low intrinsic dimensions,” which agrees with our in Table 2, at least in the large radius neighborhood. Their estimates of dimension for GPT2 is within the confidence interval we computed. They do not recognize that there might be a fiber bundle structure in play, so Lee et al. [2025] does not estimate dimensions from a smaller radius neighborhood.
Relevant to Theorem 2, there is a possibility that the implicit space of possible activations can be inferred from the behavior of the model Zhao et al. [2025b]. Since Theorem 2 represents the context window as a product of copies of the latent space, a projection of this implicit structure is the token subspace. In this vein, Zhao et al. [2025b] is not inconsistent with other findings within the literature, for instance Robinson et al. [2024]. This latter paper shows that the token subspace can be inferred (up to topological equivalence) from the responses of the LLM to single-token prompts. Taken together, Zhao et al. [2025b], Robinson et al. [2024] and the present one have pretty significant implications. Even though the token subspace may have about 0.1% of tokens that are singular, we suspect that a user will be encountering singular tokens on a regular basis when using LLMs (which seems empirically plausible). Estimating the rate of occurrence of a singular token remains an open question.
A.2 LLM Background
@@
Figure 4: Data flow in a typical LLM. A sequence of tokens forming the query is converted via the token input embedding $e^{w}$ into the initial context window, as a point in the latent space $X^{w}$. Each of these windows in the latent space are converted, token-by-token, into probability distributions via $f$ into the single token latent space $X$. From these, each token presented in the output (in the set $Y$) is obtained via a random draw. These output tokens are then used for subsequent windows.
@@
Next window prediction
@@
A.3 Computational resource requirements
Applying our entire method (starting with the token embedding matrix and ending with the $p$-values for each of the three tests at every token) to each model took approximately 12 hours of wall clock time (per model) on an Intel Core i7-3820 with 32 GB of CPU RAM and no GPU running at 3.60GHz. Almost all of the runtime was spent computing the token pairwise distance matrix $D$ in Algorithm 1. The distance matrix calculation was performed using scipy.spatial.distance_matrix(), which produces a dense matrix, by computing the distance between all pairs of tokens in the latent space. Because of the relatively low runtime requirements, we did not feel the need to use sparsity-enhancing techniques to accelerate the distance matrix calculation. The subsequent stages, including the three hypothesis tests run in only a few minutes on the same machine.
A.4 Interpretation of fiber bundles in terms of spatial variability
@@
Figure 5: Our method applied to a fiber bundle in $\mathbb{R}^{2}$. The vertical direction is the base space (long spatial scale), while the horizontal direction represents the fibers space (short spatial scale). Gray points on the right frame show estimates from a random sampling of points in the strip; the solid line shows the theoretical area versus radius curve.
@@
Figure 5 shows an example of this situation. It consists of a $1$-dimensional base space (the large radius neighborhood) and $1$-dimensional fibers (the small radius neighborhood), which in this case forms a narrow strip in the plane. Volumes (areas, in this case) of balls of small radius scale quadratically (slope $2$ in a log-log plot), but scale asymptotically linearly (slope $1$ in a log-log plot) for large radii. The transition between these two behaviors is detectable by way of a corner in the plot. This situation is easily and robustly estimated from the data; the gray points in Figure 5 (right) are derived from a random sampling of points drawn from the strip.
@@
To test for multiple spatial scales, we propose a signal model that is mathematically represented by a fiber bundle. In a fiber bundle, the long spatial scales are modeled by a space $S$, but all possible measurements are modeled by a function $p:E\to S$. The idea is that fibers $p^{-1}(b)$ are still cartesian products: pairs of both spatial scales, and these are all identical up to smooth changes of coordinates. Our method relies upon a particular geometric property of fiber bundles: we can identify if the fibers are not all identical according to when the conclusion of Theorem 1 is violated.
@@
One possible way that a fiber bundle might arise is if the underlying noise model is homoscedastic, which is to say that the dimension of the noise near a token does not depend on that token. In this case, the fibers (small radii) mostly correspond to noise, and the base (large radii) correspond mostly to signal. Nevertheless, a fiber bundle may arise with noiseless data as well; the presence of multiple spatial scales is what matters.
@@
On the other hand, if the underlying noise model is heteroscedastic—it depends on the token in question—then the fiber bundle model should be rejected. Figure 2 shows a situation that is not a fiber bundle, since there is a change in the dimension of the fiber. In the left portion of the figure, the fiber dimension is 0 while in the right portion the fiber dimension is $1$. This is detectable by looking at the volume versus radius plots for two samples. While both samples show corners in their volume versus radius plots, Theorem 1 establishes that the slopes always decrease with increasing radius for a fiber bundle. This is violated for the sample marked $\psi_{1}$, so we conclude that the space is not a fiber bundle. On the other hand, because the sample marked $\psi_{2}$ does not exhibit this violation, it is important to note that if a sample yields data consistent with Theorem 1, we cannot conclude that the space is a fiber bundle.
A.5 Mathematical proof of Theorems 1 and 2
@@
By the submersion theorem (Lee [2003]), if the Jacobian matrix of a fibered manifold $p:T\to S$ at every point has rank equal to the dimension of $S$, then the preimages $p^{-1}(\psi)\subseteq T$ of each point $\psi\in T$ are all diffeomorphic to each other. These preimages form the fibers discussed in the earlier sections of the paper.
@@
(of Theorem 1) If $e$ is assumed to be a smooth embedding [^3] Smoothness is not a strong constraint here. Any continuous function (such as ReLU) can be approximated well enough by a smooth map., the image of $e$ is a manifold of dimension $\dim T$. The pushforward of a volume form is a contravariant functor, so this means that $e_{*}v$ is the volume form for a Riemannian metric on $e(T)$. Using this Riemannian metric on $e(T)$, then [Gray, 1974, Thm 3.1] implies that for every $\psi\in e(T)$, if $r\ll\tau$, then
$(e_{*}v)\left(B_{r}(\psi)\right)=O\left(r^{\dim T}\right).$ (3)
@@
Since $e$ is assumed to be an embedding, by the tubular neighborhood theorem (Lee [2003]), it must be that $r_{2}<\tau$. Define
@@
We now address Theorem 2. Informally, we cannot get rid of all stochasticity during training (due to finite precision, stochastic gradient descent, etc.). This ensures that the transformers in practice are always in general position, hence generic.
Given a context window of w1w-1 tokens, an LLM uses its learned parameters to predict the next ($w$-th) token. This gets added to the context window and the process continues. (See Section A.2 for details.) In this section, we provide a mathematical argument showing that the use of context is insufficient to resolve singularities present in the token subspace for a generic transformer in a persistent manner. Again, denote the token subspace by $T$, noting that Section 5.2 shows that $T$ is unlikely to be a manifold for typical LLMs. Let us begin with a lemma:
Lemma 1.
@@
Consequently, the contrapositive of Lemma 1 states that if $T$ is not a manifold but $T^{w-1}$ is a manifold, it follows that $T^{w}$, the space of sequences of tokens, cannot be a manifold. At best, the latent space for every other context window size can be a manifold. For every proposed context window length, a longer window will exhibit singularities. In short, lengthening the context window does not persistently resolve singularities in the token embedding.
For instance, if we use the embedding interface provided by the LLM to embed a sequence of $w$ tokens, we will obtain a vector in the latent space $\mathbb{R}^{\ell}$. On the other hand, if we embed each of the tokens individually yet sequentially, we obtain a vector in $\mathbb{R}^{\ell\times w}$. The transformer uses the vector in $\mathbb{R}^{\ell\times w}$ to determine the vector in $\mathbb{R}^{\ell}$. Theorem 2 states that the latter ($\mathbb{R}^{\ell\times w}$) is crucial in understanding the output of the LLM. (See Section A.2 for further details.)
Proof.
(of Theorem 2) We need to manipulate the inequalities in the hypothesis to show that they satisfy [Robinson et al., 2025, Thm. 1]. Once the hypothesis is satisfied, [Robinson et al., 2025, Thm. 1] establishes that a generic set of transformers will embed $T^{w}$ into the output of the LLM. Notice that [Robinson et al., 2025, Thm. 1] considers the response of the LLM to a single token, though this can be easily extended by enlarging the bounding manifold dimension. With this in mind, the hypothesis to be satisfied is that
@@
A.6 Case study of some tokens in GPT2
Figure 6: Application of the tests to real data consisting of the tokens #, ¢, and \$ with GPT2’s token input embedding. Notice that in the case of \$, the first slope change decreases the slope; the fiber bundle test fails if the slope increases.
Figure 6 demonstrates our method on three tokens used by GPT2. The tokens \$ and # play an important role in many programming languages, whereas ¢ does not. (Note that these tokens were chosen for illustrative purposes. The $p$-value for rejecting the hypotheses is larger than $\alpha=10^{-3}$ used in Section 5.2, so these particular tokens do not appear in any of the tables in the following sections.) The token # does not show any rejections of our hypotheses. While this does not allow one to conclude that the vicinity of # is a manifold or fiber bundle, we can use Theorem 1 to estimate the dimension from the slope of the curve (approximately $53$, once logarithms of the values on both axes are taken). The token ¢ exhibits a rejection of the manifold hypothesis, but not the fiber bundle hypothesis. We can conclude that there is a irregularity at a distance of roughly $3.9$ units of ¢. The curve for \$ exhibits two slope changes at radii $3.8$ and $4.0$. The smaller radius slope change represents a violation of the manifold hypothesis, while the large radius slope change represents a violation of the fiber bundle hypothesis.
Given the presence of multiple slope changes, it is necessary to select the parameters $v_{min}$ and $v_{max}$ in Algorithm 1 carefully. Ideally, one aims to capture at most one slope change between $v_{min}$ and $v_{max}$. This is not specifically required by Algorithm 1, since it uses a sliding window, multiple rejections may be detected.
A.7 Case study of a neighborhood of ember in Mistral7B
Figure 7: This figure shows the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B. (The fiber bundle tests are not shown due to space considerations.) Since the latent space of Mistral7B has dimension $\ell=4096$, we have projected the coordinates of each token shown to $3$ using PCA. The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region. Each point shown in this figure is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test. Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature. Putative geometric structures are labeled for certain tokens.
Figures 7 and 8 show the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B. Since the latent space of Mistral7B is $4096$, we have projected the coordinates of each token shown to $3$ using PCA. The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region. Each point shown in Figure 7 is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test. Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature.
Given that the we have established that manifold rejections occur in Section 5.2, we can attempt to understand possible causes. Intuitively, several non-manifold features are visibly present [^4] The coordinates, token text, $p$-values, and a video of the space with commentary are available in the ZIP file Supplement. Due to size requirements for the coordinates of the entire vocabularies for all four LLMs, we will upload the coordinates to a public repository. in Figure 7. The tokens “ember” and “P” appear to be both putative examples of cusp points, places where the space exhibits a sharp exposed point. Manifolds cannot have cusp points, so attempting to approximate a cusp with a manifold—for instance, with UMAP—necessitates high curvature. The tokens “:” and “ly” appear to be pinch points—like what are present in a chain of sausages—which again cannot be approximated by a manifold unless it has high curvature.
@@ A.7 Case study of a neighborhood of ember in Mistral7B
Figure 8: Additional rotational views of the PCA of the neighborhood of ember in Mistral7B shown in Figure 7. Each point shown is a token. Colors are $p$-values for the manifold hypothesis, with darker colors corresponding to lower likelihood that that point’s neighborhood is a manifold with low curvature.
A.8 Table of fiber bundle rejections in all four LLMs
Table 4: Violations of the fiber bundle hypothesis
Model Token ID Token Slope change $p$-value Comment
GPT2 47623 laughable Smaller radius $9\times 10^{-6}$ Must start a word
GPT2 47128 nuance Smaller radius $2\times 10^{-4}$ Must start a word
GPT2 28664 dt Smaller radius $2\times 10^{-4}$
GPT2 37031 Mesh Smaller radius $2\times 10^{-4}$
GPT2 2689 affect Smaller radius $3\times 10^{-4}$ Must start a word
GPT2 48387 Thankfully Smaller radius $3\times 10^{-4}$
GPT2 44284 swat Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 30031 Malaysian Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 8501 Palestinian Smaller radius $7\times 10^{-4}$ Must start a word
GPT2 7864 wins Smaller radius $8\times 10^{-4}$ Must start a word
GPT2 46086 hedon Smaller radius $9\times 10^{-4}$
GPT2 17052 donor Smaller radius $9\times 10^{-4}$ Must start a word
GPT2 47482 Xan Larger radius $3\times 10^{-8}$ Must start a word
GPT2 21118 aunder Larger radius $2\times 10^{-4}$
GPT2 20564 Dri Larger radius $2\times 10^{-4}$
GPT2 1681 ney Larger radius $3\times 10^{-4}$
GPT2 2076 rodu Larger radius $3\times 10^{-4}$
GPT2 44402 Insert Larger radius $4\times 10^{-4}$
GPT2 38833 Ying Larger radius $4\times 10^{-4}$ Must start a word
Llemma7B 16352 pax Larger radius $3\times 10^{-4}$
Mistral7B 25219 änge Smaller radius $8\times 10^{-4}$
Mistral7B 9009 monitor Larger radius $8\times 10^{-5}$ Must start a word
Mistral7B 4104 H0 Larger radius $5\times 10^{-4}$
@@ A.8 Table of fiber bundle rejections in all four LLMs
A.9 Tables of manifold rejections in all four LLMs
Table 5: Rejections of the manifold hypothesis for GPT2 (1 of 2)
Number Token ID Token $p$-value Comment
1 47482 Xan 0.0000000280 Must start a word
2 28899 chemist 0.0000000602
3 32959 fry 0.00000564 Must start a word
4 47623 laughable 0.00000807 Must start a word
5 29884 Retro 0.0000108 Must start a word
6 48387 Thankfully 0.0000339
7 20619 intoler 0.0000514 Must start a word
8 18533 assumes 0.0000533 Must start a word
9 45916 handwritten 0.0000783 Must start a word
10 17615 MY 0.0000811 Must start a word
11 47128 nuance 0.000105 Must start a word
12 24817 playground 0.000110 Must start a word
13 28664 dt 0.000124
14 21118 aunder 0.000132
15 37031 Mesh 0.000135
16 37338 Disney 0.000139
17 23780 McCl 0.000155 Must start a word
18 20564 Dri 0.000196
19 17886 Destiny 0.000206 Must start a word
20 1681 ney 0.000206
21 21904 juvenile 0.000209 Must start a word
22 19980 advent 0.000211 Must start a word
23 8939 perpet 0.000213 Must start a word
24 38908 otally 0.000223
25 2689 affect 0.000226 Must start a word
26 16814 efficient 0.000238
27 2076 rodu 0.000242
28 14770 Kl 0.000260 Must start a word
29 2359 duct 0.000284
30 41890 frail 0.000292 Must start a word
31 28654 alle 0.000304 Must start a word
32 20703 Cha 0.000322 Must start a word
33 6684 $O$ 0.000333
Table 6: Rejections of the manifold hypothesis for GPT2 (2 of 2)
Number Token ID Token $p$-value Comment
34 44402 Insert 0.000336
35 38833 Ying 0.000357 Must start a word
36 41072 Georgia 0.000373
37 48679 flo 0.000384
38 32745 uddled 0.000413
39 6399 subsequ 0.000421 Must start a word
40 26106 lig 0.000439 Must start a word
41 13272 Environmental 0.000456 Must start a word
42 14840 Vladimir 0.000466 Must start a word
43 16721 Defence 0.000510 Must start a word
44 44284 swat 0.000558 Must start a word
45 28336 registering 0.000564 Must start a word
46 30031 Malaysian 0.000569 Must start a word
47 45047 trough 0.000575 Must start a word
48 11368 ears 0.000615 Must start a word
49 37070 Meh 0.000637 Must start a word
50 8501 Palestinian 0.000648 Must start a word
51 22053 reckless 0.000654 Must start a word
52 42656 Wax 0.000658 Must start a word
53 4307 Dist 0.000673 Must start a word
54 20648 Quad 0.000725 Must start a word
55 39468 relations 0.000749
56 7864 wins 0.000781 Must start a word
57 42839 KER 0.000785
58 50057 gist 0.000786 Must start a word
59 46086 hedon 0.000834
60 19055 MMA 0.000839 Must start a word
61 17052 donor 0.000865 Must start a word
62 32100 aleb 0.000873
63 18932 Asked 0.000873
64 29756 Isaiah 0.000940 Must start a word
65 13084 Survey 0.000947 Must start a word
66 40944 youngster 0.000987 Must start a word
Table 7: Rejections of the manifold hypothesis for Llemma7B
Number Token ID Token $p$-value Comment
1 24812 clojure 0.00000000434
2 10617 agu 0.000000846
3 24391 0.000000938 Cyrillic; Must start a word
4 6341 custom 0.00000163
5 26226 kW 0.00000496 Must start a word
6 19579 }ˆ{- 0.00000833
7 17485 porque 0.0000128 Must start a word
8 23146 Leopold 0.0000135 Must start a word
9 5240 Rem 0.0000232 Must start a word
10 9317 Ern 0.0000265 Must start a word
11 2466 though 0.0000540 Must start a word
12 3413 č 0.0000619 Must start a word
13 18659 endl 0.0000683 Must start a word
14 24114 corrected 0.0000738 Must start a word
15 3252 tw 0.0000788 Must start a word
16 5908 ères 0.000115
17 22385 grep 0.000151
18 2924 kind 0.000202 Must start a word
19 16352 pax 0.000243
20 8602 Tri 0.000250 Must start a word
21 18681 spielte 0.000258 Must start a word
22 22977 Heaven 0.000262 Must start a word
23 31946 0.000289 d with underline
24 7949 0.000320 Cyrillic
25 9160 Card 0.000335 Must start a word
26 14406 Wall 0.000351 Must start a word
27 29591 periodic 0.000359 Must start a word
28 10236 ucht 0.000463
29 18308 jeu 0.000553 Must start a word
30 10029 slightly 0.000807 Must start a word
31 15935 passion 0.000823 Must start a word
32 5322 Charles 0.000833 Must start a word
33 15354 victory 0.000906 Must start a word
Table 8: Rejections of the manifold hypothesis for Mistral7B
Number Token ID Token $p$-value Comment
1 20639 CV 0.000000241 Must start a word
2 1314 ember 0.00000199
3 31684 0.00000224 Korean
4 31567 ǐ 0.00000770
5 8418 rise 0.0000283 Must start a word
6 20508 poured 0.0000305 Must start a word
7 17939 aimed 0.0000364 Must start a word
8 21637 allocate 0.0000553 Must start a word
9 13215 UINT 0.0000684
10 9009 monitor 0.0000767 Must start a word
11 25867 0.000105 Cyrillic